# Enhancing Memory and Imagination Consistency in Diffusion-based World Models via Linear-Time Sequence Modeling

### v1

@@ 6.2 ViZDoom Experiments
To evaluate EDELINE’s memory retention and imagination consistency in more challenging scenarios, we conduct experiments in the first-person 3D environment ViZDoom, which presents increased complexity compared to Atari 100k’s two-dimensional perspective. Our evaluation encompasses five default scenarios (Basic, DeadlyCorridor, DefendCenter, HealthGathering, PredictPosition), with detailed scenario descriptions and reward configurations available in Appendix L. The experimental results demonstrate EDELINE’s superior performance over DIAMOND in scenarios that demand sophisticated 3D scene modeling (DeadlyCorridor) and spatiotemporal prediction (PredictPosition). To qualitatively validate the enhanced memorization and imagination capabilities of EDELINE in 3D environments, we provide qualitative visualization in Figs. 4 and 5, which illustrate EDELINE’s accurate environment detail inference in cases where DIAMOND exhibits prediction failures.
@@
Figure 5: Qualitative comparison on PredictPosition scenario. Each row shows ground truth, EDELINE’s predictions, and DIAMOND’s predictions. Numbers in boxes indicate predicted rewards, where 1 denotes a successful hit on the enemy. EDELINE correctly anticipates the reward while preserving visual fidelity.
6.3 Ablation Studies
6.3.1 Environments
The ablation studies in Sections 6.3.2 and 6.3.3 focus on five representative environments for validating the proposed key components. These include four Atari games where EDELINE demonstrates significant improvements over DIAMOND (BankHeist, DemonAttack, Hero, Seaquest), and MiniGrid-MemoryS9 for memory capability evaluation. This selection provides comprehensive validation across visual prediction quality and memorization requirements.
6.3.2 Choice of REM architecture
To validate the selection of Mamba for REM, we compare its performance against traditional linear-time sequence models GRU and LSTM across five environments. Fig. 6 illustrates that while all models achieve reasonable performance, Mamba demonstrates more stable learning curves and superior final performance, especially in memory-intensive tasks such as BankHeist and MiniGrid-MemoryS9. GRU and LSTM models exhibit increased training variance with lower final scores, which validates Mamba’s effectiveness.
Figure 8: Ablation study of harmonizer effectiveness on the Atari 100k benchmark. The bars depict mean HNS for DreamerV3 baseline (1.124), HarmonyDream (1.380, +22.8% over DreamerV3), EDELINE without harmonizers (1.674), and full EDELINE (1.866, +11.5% improvement from harmonizers). Detailed per-game scores and qualitative analysis are provided in Appendices D
6.3.3 Cross-attention in Next-Frame Predictor
To evaluate whether cross-attention improves the Next-Frame Predictor’s ability to process information-rich hidden embeddings, we examine EDELINE with and without this mechanism. As depicted in Fig. 7, the original EDELINE shows superior performance in BankHeist, MemoryS9, and Seaquest where rich contextual information processing proves essential. The MemoryS9 environment validates this necessity, as models must integrate historical information for complete representation reconstruction. Cross-attention enables effective fusion of hidden embeddings with visual features for temporal context integration.
6.3.4 Effect of Harmonizers
To validate the effectiveness of harmonizers, we evaluate performance across the full Atari 100k benchmark. The experimental results in Fig. 8 illustrate that harmonizer integration enhances DreamerV3 (HarmonyDream) performance from 1.124 to 1.380 mean HNS. Our EDELINE architecture achieves further substantial gains with harmonizer incorporation, advancing from 1.674 to 1.866 mean HNS (+11.5%). The analysis shows that while EDELINE establishes strong performance without harmonizers (1.674 mean HNS), harmonizer integration elevates performance to 1.866 mean HNS. This improvement validates the effectiveness of dynamic balance between visual and reward model learning.
@@ 7 Conclusions
In this work, we addressed the limitations of current diffusion-based world models in handling long-term dependencies and maintaining prediction consistency. Through the integration of Mamba SSMs, EDELINE effectively processed extended observation-action sequences through its recurrent embedding module, which enabled adaptive memory retention beyond fixed-context approaches. The unified framework eliminated architectural separation between observation, reward, and termination prediction, which fostered efficient representation sharing. Dynamic loss harmonization further mitigated optimization conflicts arising from multi-task learning. Extensive experiments on Atari100K, MiniGrid, and VizDoom validated EDELINE’s state-of-the-art performance, with significant improvements in quantitative metrics and qualitative prediction fidelity.
Appendix A Atari 100K training curves
@@
Appendix B Atari 100k Qualitative Analysis
To provide deeper insights into EDELINE’s superior performance, we conduct qualitative analysis on three Atari games where our approach demonstrates the most significant improvements over DIAMOND: BankHeist, DemonAttack, and Hero. Figure 10 presents temporal sequences comparing ground truth gameplay with predictions from both EDELINE and DIAMOND world models.
@@ Appendix B Atari 100k Qualitative Analysis
Appendix C Extended Performance Analysis
@@
To provide additional performance insights, we analyze EDELINE using performance profiles and optimality gaps (Agarwal et al., 2021). Fig. 11(a) shows the empirical cumulative distribution of human-normalized scores (HNS) across all 26 Atari games. The y-axis indicates the fraction of games achieving scores above each HNS threshold $\tau$. EDELINE consistently maintains a higher fraction of games across different thresholds compared to baseline methods, particularly in the middle range ($\tau$ between 1 and 4). We further analyze model performance through optimality gaps shown in Fig. 11(b). The optimality gap measures the distance between model performance and theoretical optimal behavior. EDELINE achieves the smallest optimality gap among all compared methods, demonstrating its effectiveness in approaching optimal performance across the Atari 100k benchmark suite. These detailed analyses complement the aggregate metrics presented in Section 6.1 by providing a more granular view of performance distribution and theoretical efficiency.
@@
Figure 12: Comparison of state-of-the-art model-based RL methods without using look-ahead search techniques on the Atari 100k benchmark. EDELINE outperforms all existing model-based approaches on the Atari 100k benchmark. Previous methods can be categorized by their world model architectures: Transformer-based models (TWM (Robine et al., 2023), IRIS (Micheli et al., 2023), REM (Cohen et al., 2024), STORM (Zhang et al., 2023), $\Delta$-IRIS (Alonso et al., 2023), TWISTER (Anonymous, 2025)), RNN-based models (DreamerV3 (Hafner et al., 2024), PaMoRL (Wang et al., 2024), HarmonyDream (Ma et al., 2024)), SSM-based models (Drama (Wenlong et al., 2024)), and Diffusion-based models (DIAMOND (Alonso et al., 2024)). Our proposed EDELINE advances the state-of-the-art by integrating diffusion modeling with state space models, combining their respective strengths in visual generation and temporal modeling.
Appendix D Harmonizers Ablation Results

### v2

@@ 6.3 ViZDoom Experiments
Figure 4: Training curves comparing EDELINE (blue) and DIAMOND (green) across five ViZDoom scenarios. The solid lines represent the mean performance over three seeds, with shaded regions indicating standard deviation.
To evaluate EDELINE’s memory retention and imagination consistency in more challenging scenarios, we conduct experiments in the first-person 3D environment ViZDoom, which presents increased complexity compared to Atari 100k’s two-dimensional perspective. Our evaluation encompasses five default scenarios (i.e., Basic, DeadlyCorridor, DefendCenter, HealthGathering, PredictPosition), with detailed scenario descriptions and reward configurations available in Appendix D.10. The experimental results presented in Fig. 4 demonstrate EDELINE’s superior performance over DIAMOND in scenarios that demand sophisticated 3D scene modeling (DeadlyCorridor) and spatiotemporal prediction (PredictPosition).
The DeadlyCorridor scenario provides a particularly compelling demonstration of EDELINE’s memory advantages. This environment requires the agent to navigate a corridor, eliminate enemies on both sides, and ultimately reach armor at the corridor’s end. The task necessitates sustained spatial awareness throughout a relatively long trajectory. As Fig. 5 illustrates, EDELINE’s SSM effectively integrates information from the entire history, whereas DIAMOND remains constrained to only the past four frames. This memory limitation significantly impairs DIAMOND’s ability to maintain consistent predictions in 3D environments where current observations provide only partial information about the scene. The visual comparison reveals that EDELINE accurately models both the game environment and particle effects while correctly predicts the corridor’s end and armor object in the final steps. This substantiates EDELINE’s comprehension of the agent’s position relative to the corridor’s end through its extended memory capabilities, while DIAMOND’s predictions become increasingly blurry and inconsistent due to its memory constraints. It also reveals that the enhanced spatial awareness directly contributes to EDELINE’s superior performance in the environment.
@@ 7 Conclusions
In this work, we addressed the limitations of current diffusion-based world models in handling long-term dependencies and maintaining prediction consistency. Through the integration of Mamba SSMs, EDELINE effectively processed extended observation-action sequences through its recurrent embedding module, which enabled adaptive memory retention beyond fixed-context approaches. The unified framework eliminated architectural separation between observation, reward, and termination prediction, which fostered efficient representation sharing. Dynamic loss harmonization further mitigated optimization conflicts arising from multi-task learning. Extensive experiments on Atari 100K, Crafter, MiniGrid, and VizDoom validated EDELINE’s state-of-the-art performance, with significant improvements in quantitative metrics and qualitative prediction fidelity.
Appendix A Limitation
Although EDELINE maintains computational efficiency equivalent to that of DIAMOND and simultaneously addresses its memory limitations, the fundamental computational requirements of diffusion-based world models remain substantial relative to alternative approaches. The iterative characteristics of the diffusion process necessitate multiple forward passes throughout both training and inference phases, which consequently elevates computational overhead. The enhancement of computational efficiency within diffusion processes for world modeling applications constitutes a promising avenue for future research endeavors.
Appendix B Boarder Impact
EDELINE addresses fundamental limitations inherent in diffusion-based world models through the enhancement of memory capabilities and imagination consistency. Our research contributes three pivotal advances to data-efficient reinforcement learning: (1) we overcome the fixed context window constraints of prior methodologies such as DIAMOND through the implementation of selective state space models, which substantially enhances memory capabilities, (2) we establish an end-to-end training methodology that effectively leverages hidden embeddings and incorporates HarmonyDream to augment world modeling performance, and (3) we demonstrate superior performance across visually rich 2D environments (i.e., Atari), challenging 3D first-person perspectives (i.e., ViZDoom), and memory-intensive procedurally-generated survival scenarios (i.e., Crafter), which illustrates the potential of our approach to address increasingly complex domains.
World models represent a promising direction for the improvement of sample efficiency and safety in reinforcement learning through the reduction of direct environment interaction requirements. While EDELINE advances this research domain, we acknowledge that imperfections in world models may continue to result in suboptimal agent behaviors. We believe our contributions toward more accurate and memory-capable world models will serve to mitigate such risks in future applications.
Appendix C Additional Background Material Section
C.1 Score-based Diffusion Generative Models
Diffusion probabilistic modeling [59, 19, 60] and score-based generative modeling [21, 61, 62] can be unified through a forward stochastic differential equation (SDE) formulation [22]. The forward diffusion process $\{{\mathbf{x}}^{\tau}\}$ with continuous time variable $\tau$ transforms the data distribution $p^{0}=p^{\text{data}}$ to prior distribution $p^{T}=p^{\text{prior}}$, expressed as:
where ${\mathbf{n}}$ is Gaussian noise with zero mean and variance determined by a variance scheduler $\sigma(\tau)$ that follows a noise distribution $p^{\text{train}}$, and $({\mathbf{x}}^{0}+{\mathbf{n}})$ corresponds to the perturbed data ${\mathbf{x}}^{\tau}$. The score function can then be estimated through: $\nabla_{{\mathbf{x}}}\log p^{\tau}({\mathbf{x}})=\frac{1}{\sigma^{2}}(D_{\theta}({\mathbf{x}};\sigma)-{\mathbf{x}})$. In practice, modeling the denoiser $D_{\theta}$ directly can be challenging due to the wide range of noise scales. To address this, EDM [63] introduces a design space that isolates key design choices, including preconditioning functions $\{c_{\text{skip}},c_{\text{out}},c_{\text{in}},c_{\text{noise}}\}$ to modulate the unconditioned neural network $F_{\theta}$ to represent $D_{\theta}$, which can be formulated as:
C.2 Multi-task Essence of World Model Learning
Modern world models [15, 7, 2, 8] typically address two fundamental prediction tasks: the modeling of environment dynamics through observations and the prediction of reward signals. The learning of these tasks requires distinct considerations based on the complexity of the environment. In simple low-dimensional settings, separate learning approaches suffice for each task. However, the introduction of high-dimensional visual inputs fundamentally alters this paradigm, as partial observability creates an inherent coupling between state estimation and reward prediction. This coupling necessitates joint learning through shared representations, an approach that aligns with established multi-task learning principles [64]. The implementation of such joint learning through shared representations introduces several technical challenges. The integration of multiple learning objectives requires careful consideration of their relative importance and interactions. A fundamental difficulty stems from the inherent scale disparity between high-dimensional visual observations and scalar reward signals. This disparity manifests in the world model learning objective, which combines observation modeling $\mathcal{L}_{o}(\theta)$, reward modeling $\mathcal{L}_{r}(\theta)$, and dynamics modeling $\mathcal{L}_{d}(\theta)$ losses with weights $w_{o}$, $w_{r}$, $w_{d}$ to control relative contributions:
HarmonyDream [55] demonstrated that observation modeling tends to dominate this objective due to visual inputs’ high dimensionality compared to scalar rewards. Their work introduced a variational formulation:
$\mathcal{L}(\theta,w_{o},w_{r},w_{d})$ $=\sum_{i\in\{o,r,d\}}\mathcal{H}(\mathcal{L}_{i}(\theta),\frac{1}{w_{i}})=\sum_{i\in\{o,r,d\}}w_{i}\mathcal{L}_{i}(\theta)+\log(\frac{1}{w_{i}}),$ (12)
Appendix D Additional Experiments and Experiment Details
In this section, we provide extensive additional experiments and detailed experimental setups that complement the primary results presented in the main manuscript. These supplementary materials include in-depth analyses, ablation studies, and technical details that further validate and illuminate EDELINE’s advantages.
D.1 Experimental Setup
We evaluate EDELINE on the Atari 100k benchmark [15], which serves as the standard evaluation protocol in recent model-based RL literature for fair comparison. In addition, our experimental validation extends to ViZDoom [18], MiniGrid [16], and Crafter [17] environments to demonstrate broader applicability. To ensure statistical significance, all reported results represent averages across three independent runs.
We evaluate EDELINE on the Atari 100k benchmark [15], which serves as the standard evaluation protocol in recent model-based RL literature for fair comparison. In addition, our experimental validation extends to ViZDoom [18] and MiniGrid [16] environments to demonstrate broader applicability. To ensure statistical significance, all reported results represent averages across three independent runs. The Atari 100k benchmark [15] encompasses 26 diverse Atari games that evaluate various aspects of agent capabilities. Each agent receives a strict limitation of 100k environment interactions for learning, in contrast to conventional Atari agents that typically require 50 million steps. EDELINE’s performance is evaluated against current state-of-the-art world model-based approaches, including DIAMOND [8], STORM [33], DreamerV3 [2], IRIS [35], TWM [32], and Drama [48]. For evaluating 3D scene understanding capabilities, we employ VizDoom scenarios that demand sophisticated 3D spatial reasoning in first-person environments. This provides a crucial testing ground beyond the third-person perspective of Atari environments. Furthermore, the Crafter and MiniGrid memory scenarios evaluate memorization capabilities through tasks that require information retention across extended time horizons.
D.2 EDM Network Preconditioners and Training
Following DIAMOND [8], we use the EDM preconditioners from [63] for normalization and rescaling to improve network training:
$c_{in}^{\tau}=\frac{1}{\sqrt{\sigma(\tau)^{2}+\sigma_{data}^{2}}}$ (13)
$c_{noise}^{\tau}=\frac{1}{4}\log(\sigma(\tau))$ (15)
where $\sigma_{data}=0.5$.
$\log(\sigma(\tau))\sim\mathcal{N}(P_{mean},P_{std}^{2}),$ (17)
where $P_{mean}=-0.4$, $P_{std}=1.2$.
D.3 Atari 100K Training Curves
Fig. 6 presents the detailed training curves for all individual games in the Atari100k benchmark.
@@ D.3 Atari 100K Training Curves
D.4 Atari 100k Qualitative Analysis
To provide deeper insights into EDELINE’s superior performance, we conduct qualitative analysis on three Atari games where our approach demonstrates the most significant improvements over DIAMOND: BankHeist, DemonAttack, and Hero. Fig. 7 presents temporal sequences comparing ground truth gameplay with predictions from both EDELINE and DIAMOND world models.
@@ D.4 Atari 100k Qualitative Analysis
D.5 Extended Atari 100k Benchmark Performance Analysis
@@
To provide additional performance insights, we analyze EDELINE using performance profiles and optimality gaps [57]. Fig. 8(a) shows the empirical cumulative distribution of human-normalized scores (HNS) across all 26 Atari games. The y-axis indicates the fraction of games achieving scores above each HNS threshold $\tau$. EDELINE consistently maintains a higher fraction of games across different thresholds compared to baseline methods, particularly in the middle range ($\tau$ between 1 and 4). We further analyze model performance through optimality gaps shown in Fig. 8(b). The optimality gap measures the distance between model performance and theoretical optimal behavior. EDELINE achieves the smallest optimality gap among all compared methods, demonstrating its effectiveness in approaching optimal performance across the Atari 100k benchmark suite. These detailed analyses complement the aggregate metrics presented in Section 6.1 by providing a more granular view of performance distribution and theoretical efficiency.
@@
Figure 9: Comparison of state-of-the-art model-based RL methods without using look-ahead search techniques on the Atari 100k benchmark. EDELINE outperforms all existing model-based approaches on the Atari 100k benchmark. Previous methods can be categorized by their world model architectures: Transformer-based models (TWM [32], IRIS [35], REM [65], STORM [33], $\Delta$-IRIS [58], TWISTER [66]), RNN-based models (DreamerV3 [2], PaMoRL [67], HarmonyDream [55]), SSM-based models (Drama [48]), and Diffusion-based models (DIAMOND [8]). Our proposed EDELINE advances the state-of-the-art by integrating diffusion modeling with state space models, combining their respective strengths in visual generation and temporal modeling.
D.6 Atari 100k Linear Probing Analysis
To quantitatively evaluate the information content captured in EDELINE’s hidden representations, we conducted a comprehensive linear probing analysis, which assesses how effectively the model’s hidden states encode information critical for both reward prediction and observation prediction tasks.
Table 3: Linear probing loss of reward and observation prediction from EDELINE and DIAMOND hidden states. $\mathcal{L}_{\text{rew}}$ (reward loss) was measured using cross-entropy loss on reward prediction, while $\mathcal{L}_{\text{obs}}$ (observation loss) was measured using MSE loss on predicting encoded observation features. EDELINE outperforms DIAMOND in observation prediction with an average 57.3% reduction in loss, while both models achieve comparable performance on reward prediction (i.e., average $\mathcal{L}_{\text{rew}}$ of 0.0246 vs. 0.0258). These results demonstrate that EDELINE’s unified hidden representation successfully encodes information required for both tasks.
Environment EDELINE $\mathcal{L}_{\text{rew}}$ DIAMOND $\mathcal{L}_{\text{rew}}$ EDELINE $\mathcal{L}_{\text{obs}}$ DIAMOND $\mathcal{L}_{\text{obs}}$
Alien 0.0434 0.0663 0.6161 1.1798
Amidar 0.0186 0.0187 0.6430 0.7520
Assault 0.0779 0.0270 0.4894 0.4199
Asterix 0.0114 0.0018 0.8247 1.4411
BankHeist 0.0023 0.0039 0.0312 2.5398
BattleZone 0.0129 0.0220 0.2265 0.1413
Boxing 0.0482 0.0011 0.3089 0.5328
Breakout 0.0615 0.0264 0.6224 0.4871
ChopperCommand 0.0261 0.0113 0.6815 1.3819
CrazyClimber 0.0029 0.0034 0.2844 0.5513
DemonAttack 0.0609 0.0726 0.8588 1.0039
Freeway 0.0000 0.0000 0.2163 0.3911
Frostbite 0.0001 0.0002 0.1992 0.7435
Gopher 0.0090 0.0091 0.3139 0.9558
Hero 0.0065 0.0088 0.0708 0.1013
Jamesbond 0.0054 0.0058 0.5931 1.1080
Kangaroo 0.0009 0.0046 0.0992 0.5054
Krull 0.0728 0.0455 0.3216 0.8615
KungFuMaster 0.0048 0.0043 0.2674 0.8735
MsPacman 0.0044 0.0669 0.2510 0.8769
Pong 0.0000 0.0000 0.1770 1.3162
PrivateEye 0.0014 0.0144 0.1188 0.4931
Qbert 0.0240 0.0025 0.1813 0.2530
RoadRunner 0.0658 0.0310 0.2032 0.7896
Seaquest 0.0008 0.0299 0.2159 0.7852
UpNDown 0.1012 0.2110 0.5967 1.0783
ViZDoom-DeadlyCorridor 0.0008 0.0094 0.3354 1.2848
Average Loss 0.0246 0.0258 0.3610 0.8462
We trained linear probes on the hidden states extracted from both EDELINE and DIAMOND models. For our experimental setup, we collected 50 trajectories as training data and 10 trajectories as validation data for each environment, with five independent random seeds employed to ensure statistical robustness. For reward prediction, we trained a single linear layer to predict rewards for 50 epochs. For observation feature prediction, we trained a multi-layer perceptron (MLP) to predict encoded features of the subsequent observation (extracted from the pre-trained actor-critic network encoder) for 100 epochs. For EDELINE, we utilized the Mamba hidden states, while for DIAMOND, we employed the LSTM hidden states from its reward-termination model. We conducted these experiments across 26 Atari environments and the challenging ViZDoom-DeadlyCorridor environment to ensure comprehensive evaluation.
Table 3 presents the results of our linear probing analysis. EDELINE and DIAMOND demonstrate comparable performance on reward prediction, with average losses of 0.0246 and 0.0258 respectively. EDELINE outperforms DIAMOND in 15 out of 27 environments on this task. More significantly, EDELINE substantially outperforms DIAMOND in observation prediction, with an average $57.3\%$ reduction in loss. EDELINE achieves lower observation prediction loss in 24 out of 27 environments. The substantial improvement in observation prediction while maintaining comparable reward prediction performance demonstrates that EDELINE’s unified hidden representation successfully captures information required for both tasks.
D.7 Generation Quality Evaluation
To quantitatively assess the generation quality of our proposed world model, we conducted a comprehensive evaluation measuring pixel-wise Mean Squared Error (MSE) across 26 Atari environments and the challenging ViZDoom-DeadlyCorridor environment. This analysis provides direct evidence of EDELINE’s improved predictive accuracy compared to DIAMOND.
Table 4: Pixel-wise MSE comparison between EDELINE and DIAMOND across 27 environments. Lower values (in bold) indicate better performance. The rightmost column shows the imagination horizon length for each environment. The bottom row reports the average normalized score (EDELINE MSE / DIAMOND MSE), with values below 1.0 indicating EDELINE’s overall superior performance.
Environment EDELINE DIAMOND Imagine Horizon Length
Alien 0.0063 0.0064 635
Amidar 0.0177 0.0235 1029
Assault 0.1011 0.2528 566
Asterix 0.0190 0.0177 1493
BankHeist 0.1510 0.1733 2019
BattleZone 0.0306 0.0292 1407
Boxing 0.0068 0.0174 1371
Breakout 0.0420 0.0415 1796
ChopperCommand 0.0084 0.0082 3006
CrazyClimber 0.0666 0.0686 3219
DemonAttack 0.0312 0.0318 1474
Freeway 0.0015 0.0030 1986
Frostbite 0.0338 0.0370 564
Gopher 0.0152 0.0172 2689
Hero 0.1559 0.2005 2103
Jamesbond 0.2967 0.3023 2212
Kangaroo 0.0061 0.0063 3241
Krull 0.1577 0.1575 1326
KungFuMaster 0.0210 0.0210 1864
MsPacman 0.0179 0.0197 892
Pong 0.0035 0.0034 1532
PrivateEye 0.1043 0.0638 2454
Qbert 0.0435 0.0460 1434
RoadRunner 0.0526 0.0532 960
Seaquest 0.0107 0.0136 1810
UpNDown 0.1223 0.1234 1279
ViZDoom-DeadlyCorridor 0.0179 0.0184 73
Mean Normalized Score 0.918 1.000
Table 4 presents the comprehensive results of our pixel-wise MSE comparison. The analysis reveals that EDELINE achieves lower MSE than DIAMOND in 20 out of 27 environments. For meaningful comparison across environments with varying visual complexity, we calculated the average normalized score (EDELINE MSE / DIAMOND MSE) across all environments. Our analysis demonstrates that EDELINE achieves an average normalized score of 0.918, which represents an overall $8.2\%$ reduction in pixel-level error compared to DIAMOND.
The superior generation quality of EDELINE can be attributed to its utilization of Mamba to incorporate longer observation history beyond the four frames employed by DIAMOND. This capability enables EDELINE to maintain more consistent predictions over extended horizons, which proves particularly advantageous in environments that necessitate memory (such as ViZDoom) or involve complex dynamics.
D.8 Training Time Profile
We performed detailed training time profiling of EDELINE to evaluate its computational efficiency compared to DIAMOND. Table 5 provides a comprehensive breakdown of training time components across different scales, while Table 6 directly compares EDELINE with DIAMOND. Our profiling reveals that EDELINE achieves notable efficiency improvements in world model training. For world model updates, EDELINE is approximately 26.8% faster than DIAMOND. This efficiency gain stems from our unified architecture approach. While DIAMOND employs a two-stage training process (diffusion model for observations plus a separate CNN-LSTM network for reward/termination prediction), EDELINE’s unified architecture enables joint learning of observations, rewards, and terminations through shared representations. In addition, Mamba’s parallel scan algorithm contributes to this computational advantage during training. For actor-critic updates, EDELINE requires approximately $17.7\%$ more computation time than DIAMOND. This difference occurs because while both methods use identical actor-critic architectures, EDELINE’s world model involves more complex inference during imagined rollouts due to SSM processing and cross-attention mechanisms. Specifically, as shown in Table 5, each imagination step requires 24.4ms, with 17.2ms dedicated to observation prediction and 6.5ms to reward/termination prediction. Overall, these timing differences balance out, resulting in comparable total training time between the two approaches. This demonstrates that EDELINE successfully maintains computational efficiency comparable to DIAMOND while delivering significantly enhanced memory capabilities and performance.
Single update Time (ms) Detail (ms)
Total 548.6 148.6 + 400
   World model update 148.6 -
   Actor-Critic model update 400 15 $\times$ 24.4 + 34
    Imagination step (x 15) 24.4 17.2 + 6.5 + 0.7
     Next observation prediction 17.2 -
     Denoising step (x 3) 5.7 -
     Reward/Termination prediction 6.5 -
     Action prediction 0.7 -
    Loss computation and backward 34 -
Epoch Time (s) Detail (s)
Total 219.4 59.4 + 160
   World model 59.4 400 $\times$ 148.6 ×$\times 10^{-3}$
   Actor-Critic model 160 400 $\times$ 400 ×$\times 10^{-3}$
Run Time (days) Detail (days)
Total 2.9 2.5 + 0.4
   Training time 2.5 1000 $\times$ 219.4 / (24 $\times$ 3600)
   Other (collection, evaluation, checkpointing) 0.4 -
Module DIAMOND (ms/s/days) EDELINE (ms/s/days) Difference (%)
Single Update
Total 543 ms 548.6 ms +1.03%
   World Model Update 203 ms 148.6 ms -26.80%
   Actor-Critic Model Update 340 ms 400 ms +17.65%
Epoch
Total 217 s 219.4 s +1.11%
   World Model 81 s 59.4 s -26.67%
   Actor-Critic Model 136 s 160 s +17.65%
D.9 Ablation Studies
To systematically validate the effectiveness of EDELINE’s key architectural components, we perform comprehensive ablation studies across multiple environments. These studies isolate the contribution of each design decision and demonstrate their impact on model performance. The ablation studies in Appendix D.9.1 and D.9.2 focus on five representative environments for validating the proposed key components. These include four Atari games where EDELINE demonstrates significant improvements over DIAMOND (BankHeist, DemonAttack, Hero, Seaquest), and MiniGrid-MemoryS9 for memory capability evaluation. This selection provides comprehensive validation across visual prediction quality and memorization requirements. To further validate the advantages of MAMBA compared to Transformer architectures with quadratic time complexity, we conducted experiments on memory tasks such as MiniGrid MemoryS7/S9 and Crafter. These results are presented in Appendix D.9.4.
D.9.1 Choice of REM architecture
To validate the selection of Mamba for REM, we compare its performance against traditional linear-time sequence models GRU and LSTM across five environments. Fig. 10 illustrates that although all models achieve reasonable performance, Mamba demonstrates more stable learning curves and superior final performance, particularly in memory-intensive tasks such as BankHeist and MiniGrid-MemoryS9. GRU and LSTM models exhibit increased training variance with lower final scores, which validates Mamba’s effectiveness.
D.9.2 Cross-Attention in Next-Frame Predictor
To evaluate whether cross-attention improves the Next-Frame Predictor’s ability to process information-rich hidden embeddings, we examine EDELINE with and without this mechanism. As depicted in Fig. 11, the original EDELINE demonstrates superior performance in BankHeist, MemoryS9, and Seaquest where rich contextual information processing proves essential. The MemoryS9 environment validates this necessity, as models must integrate historical information for complete representation reconstruction. Cross-attention enables effective fusion of hidden embeddings with visual features for temporal context integration.
D.9.3 Effect of Harmonizers
To validate the effectiveness of harmonizers, we evaluate performance across the full Atari 100k benchmark.